. At the end of 2017, Company Xs P/B multiple was closest to: 7.8 8.0 8.3 9.3 10.0 5. At the end of 2017, Company Xs reported marketable securities closest to: $90 $50 $40 $30 $20 16. At the end of 2017, Company X reported EBITDA closest to: $430 $420 $360 $320 e. $310 17.At the end of 2017, Company Xs effective tax was closest to: a. 20% b. 23% c. 25% d. 27% e, 33% 18. Do not include marketable securities or the current portion of long term debt when calculating working capital investment. In 2018, Company Xs Cash Flow from Operations was closest to a. $180 b. $220 c. $230 d. $280 e. $290 19. Ignore your answer to question 18 and assume that CFO was $250. Calculate fixed capital investment as the year on year change in Gross PPE. In 2018, Company Xs Free cash Flow to Firm was closest to a. $110 b. $150 c. $170 d. $190 e. $210 20. In 2018, Company Xs Return on Equity, where ROEENI/EOY SE, is closest to: 20% 23% a. b. c. d. e. 25% 27% 33%Use the following information about Company X to help answer questions 14-20 Company X went public at the start of 2017. At the time of the IPO the company: Issued 200M shares of stock at $25 per share (thus raising $500M in equity) issued $400M of long term debt; as part of this interest only loan, the company agreed to pay creditors 15.0% per year; this implies the current portion of long term debt is SO in 201 7/2018 · At the end of 2017, Company X issued financial statements. A complete summary is below: $400M in sales, $260M in operating earnings, and $160 in net income .$500M Gross PPE , $450 Net PPE . Working Capital of $430, excluding marketable securities The company paid a dividend of 30 cents per share Investors were pleased with the results; Company Xs share price closed at $28 per share At the end of 2018, Company X issued financial statements. A complete summary is below $480M in sales, $300M in operating earnings, and $200 in net income $600M Gross PPE, $490 Net PPE Working Capital of $460, excluding marketable securities The company paid a dividend of 40 cents per share Investors were pleased with the results; Company Xs share price closed at $33 per share . Throughout 2017 and 2018 The company did not issue or repurchase any shares of stock other than during the IPO . The company did not issue or retire any long term debt other than during the initial deht aff
Answer #1

Please note that there appears to be a small typographical error in the statement:

"...Issued 200M shares of stock at at $ 25 per share thus raising $ 500 M in equity"

If we multiply, nos. of shares issued with the issue price, we get the quantum of equity raised as = 200M x 25 = $ 5,000 M and not $ 500 M as stated above. In order to rectify this I am assuming the issue price was $ 2.5 / share and not $ 25 / share.

Question 14

Price at the end of 2017, P = $ 28

Nos. of shares outstanding = nos. of shares issued at the time of IPO = N = 200 M

Hence, market capitalization = P x N = 28 x 200 = $ 5,600 M

Book value at the end of 2017 = Book value at the beginning of 2017 + Net income during 2017 - dividend payout

Book value at the beginning of 2017 = equity raised at the time of IPO = $ 500 M

Net income during 2017 = $ 160 M

Dividend paid during 2017 = dividend per share x shares outstanding = 0.30 x 200 = $ 60 M

Hence, Book value at the end of 2017 = Book value at the beginning of 2017 + Net income during 2017 - dividend payout = 500 + 160 -60 = $ 600 M

Hence price to book ratio = P / B = Market capitalization / Book value of equity = 5,600 / 600 = 9.3

Hence, correct answer is option (d) 9.3

------------------------

Question 15

On the asset side, the company will have following:

  • Net PPE = 450
  • working capital excluding marketable securities = 430
  • Marketable securities

On the liability and equity side, it will have:

  • Book value of debt = 400
  • Book value of equity = 600 (calculated in Question 14 above)

Recall basic equation of accounting:

Assets = Liabilities + Equity

Hence, Net PPE + Working capital excluding marketable securities + marketable securities = Book value of debt + book value of equity

Hence, 450 + 430 + marketable securities = 400 + 600 = 1,000

Hence, marketable securities = 1,000 - 450 - 430 = 120

None of the options offer 120 as a choice. Hence, there appears to be an error. None of the options is correct.

--------------------------------

Question 16

Operating income in 2017 = 260 = EBIT

Gross PPE = 500

Net PPE = 450 = Gross PPE - accumulated Depreciation till 2017

Since 2017 is the first year of operation, accumulated depreciation till 2017 = depreciation during 2017

Hence, Net PPE = 450 = Gross PPE - accumulated Depreciation till 2017 = Gross PPE - depreciation during 2017 = 500 - depreciation during he year.

Hence, depreciation during the year 2017 = 500 - 450 = 50

Hence, EBITDA = EBIT + Depreciation = 260 + 50 = 310.

Hence, correct answer is option (e) $ 310

----------------------

Question 17

EBIT = 260

Interest = 15% of loan outstanding = 15% x 400 = 60

Hence, EBT = EBIT - Interest = 260 - 60 = 200

Net income = 160 = EBT - taxes = 200 - taxes

Hence, taxes = 200 - 160 = 40

Hence, effective tax rate = Taxes / EBT = 40 / 200 = 0.2 = 20%

Hence, correct answer is option (a) 20%

-------------------------

Question 18

Cash flow from operations = Net Income + Interest + Non cash expenses - Increase in working capital

Net income for 2018 = 200

Interest = 15% of debt = 15% x 400 = 60

Non cash expense = Depreciation expense for the year 2018

Net PPE at the end of 2018 = Net PPE at the end of 2017 + assets purchased during 2018 - depreciation during 2018

Asset purchased during 2018 = Gross PPE at the end of 2018 - gross PPE at the end of 2017 = 600 - 500 = 100

Hence,  Net PPE at the end of 2018 = Net PPE at the end of 2017 + assets purchased during 2018 - depreciation during 2018

Hence, 490 = 450 + 100 - depreciation during 2018

Hence, depreciation during 2018 = 450 + 100 - 490 = 60

Increase in working capital during 2018 = Working capital at the end of 2018 - working capital at the end of 2017 = 460 - 430 = 30

Cash flow from operations = Net Income + Interest + Non cash expenses - Increase in working capital = 200 + 60 + 60 - 30 = 290

Hence, correct answer is option (e) 290

---------------------------

Question 19

Cash flow from operations, CFO = 250

Capital expenditure during 2018 = Asset purchased during the year 2018 = 100 (calculated in Question 18 above)

Hence, Free cash flow tot he firm = CFO - Capital expenditure = 250 - 100 = 150

hence, correct answer is option (b) $ 150

------------

Question 20

ROE = net Income / End of year shareholders' equity

Net income = 200

End of year shareholders' equity = Shareholder's equity at the end of year 2017 + net income during 2018 - dividends paid during 2018 = 600 (as calculated in Question 14) + 200 - dividend of 40 cents per share x nos. of shares = 600 + 200 - 0.40 x 200 = 720

Hence, ROE = 200 / 720 = 0.2778 = 27.78%

hence correct answer is option (d) 27%